x to the power of 2 = 1/4 and x to the power of 2 = 1/9 plzzz help i dont understand and dont get it

Answer by checkley71(8403)

X^2=1/4
X=SQRT1/4
X=1/2 ANSWER.
PROOF
1.2*1/2=1/4
1/4=1/4
---------------------------
X^2=1/9
X=SQRT1/9
X=1/3 ANSWER.
PROOF
1/3*1/3=1/9
1/9=1/9